OpenStudy (anonymous):

A die is rolled 20 times and the number of twos that come up are tallied. find the propabilty of getting exactly four twos

OpenStudy (anonymous):

A. 0.202 B. 0.075 C. 0.101 D. 0.083

OpenStudy (anonymous):

which one would it be?

OpenStudy (anonymous):

that was wrong it would be \[\dbinom{20}{4}(\frac{1}{6})^4(\frac{5}{6})^{16}\]

OpenStudy (anonymous):

reasoning that you got 4 two, each of which has probability \(\frac{1}{6}\) and therefore 14 not two each with probability \(\frac{5}{6}\) and \(\binom{20}{4}\) different was to get that

OpenStudy (anonymous):

i keep thinking 14 but of course 20 - 4 = 16

OpenStudy (anonymous):

but i needed it round it to a deciaml yaa now(:

OpenStudy (anonymous):

.2022 it looks like

OpenStudy (anonymous):

go with answer 1

OpenStudy (anonymous):

yes(: haaaa thanks so muuch(: . wanna help me with another ?

OpenStudy (anonymous):

on line class?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

yeaaah lol

OpenStudy (anonymous):

Two marbles are drawn without replacement from a box with 3 white, 2 green, 2 red, and 1 blue marble. Find the probability. The second marble is white given the first marble is blue.

OpenStudy (anonymous):

A. B. C. D.

OpenStudy (anonymous):

it has to be a fraction

OpenStudy (anonymous):

oh this we just think one marble is gone, it was a blue one there are now 7 marbles in the bag, 3 are white \[\frac{3}{7}\]

OpenStudy (anonymous):

hope that is an answer choice, and it was much easier to compute than the last one

OpenStudy (anonymous):

thanks so much(: i hate all this math bullpellet its hardd .

OpenStudy (anonymous):

yeah probability can be a pain but you have to admit the last one didn't require much right? 7 marbles left, 3 are white? not like the last one

OpenStudy (anonymous):

haa yeaah way less harder(: